# Rokhlin dimension for actions of residually compact groups

### v1

@@ 1. Introduction
The Rokhlin dimension was initially defined for actions by finite groups and the group $\mathbb{Z}$ of integers [HWZ15] and subsequently generalized to actions of more general groups: compact groups in [Gar17b, Gar17a, GHS21], residually finite groups in [SWZ19], and the group $\mathbb{R}$ of real numbers in [HSWW17]. All of these definitions can be used to transfer good structure properties, such as finite nuclear dimension and $\mathcal{Z}$-absorption (the latter case requires an extra assumption of commuting towers), from a C*-algebra $A$ to its crossed product C*-algebra $A\rtimes_{\alpha}G$.
In this paper, we unify all the known constructions of Rokhlin dimension and generalize the theory to work for residually compact groups, a concept that extends residual finiteness to the scope of locally compact groups. A second countable locally compact group is called residually compact if it admits a sequence consisting of closed cocompact subgroups such that the intersection of these subgroups is trivial. Such a sequence is called a residually compact approximation. The definition of residually compact groups in this paper is slightly weaker than that of Amini in [Ami24], in that we do not need the sequence to be decreasing. On the other hand, just like the case of residually finite groups, in order to apply the theory of Rokhlin dimension, we shall almost always assume a regularity condition on the residually compact approximations. Note that the various classes of groups mentioned in the last paragraph, namely compact groups, residually finite groups and the group $\mathbb{R}$ of real numbers, are all residually compact groups. Further natural examples, for which a Rokhlin dimension theory has not been established, include $\mathbb{R}^{d}$ and the continuous Heisenberg group, etc. In fact, all these groups admit regular residually compact approximations.
@@
To see how this definition generalizes the description given earlier that uses Rokhlin towers, we observe that each map $\varphi_{l}$ in the above definition can be lifted to an approximately equivariant, approximately central, approximately order zero, completely positive contractive map $\psi_{l}\colon C(G/H)\to A$ (see Proposition 5.3). In particular, if $G$ is a finite group and $H$ is chosen to be trivial, then $\{\psi_{l}(\delta_{g})\colon g\in G\}$ forms a Rokhlin tower in the previous descriptions, where each $\delta_{g}$ is the characteristic function of $\{g\}$.
With this definition, we are able to unify and generalize the previous theories on Rokhlin dimension. In particular, we show a number of important C*-algebraic properties are preserved under taking crossed products with regard to actions with finite Rokhlin dimension. Let us start with finite nuclear dimension.
@@
Another regularity property of great significance in the modern theory of C*-algebras is $\mathcal{Z}$-stability, also known as $\mathcal{Z}$-absorption, that is, the ability to tensorially absorb the Jiang-Su algebra $\mathcal{Z}$. This fits into a range of interesting properties called $D$-absorption, where $D$ is a strongly self-absorbing C*-algebra, e.g., $\mathcal{Z}$, $M_{p^{\infty}}$, $\mathcal{O}_{\infty}$ and $\mathcal{O}_{2}$ ([TW07]). For example, $\mathcal{O}_{\infty}$-absorption has played a pivotal role in the classification theory of purely infinite C*-algebras. Again, we extend the previous theories by showing that, under the assumptions of finite Rokhlin dimension with commuting towers and finite-dimensional box spaces, $D$-absorption is also preserved under taking crossed products.
@@ 5. Rokhlin dimension for actions of residually compact groups
In this section, we introduce the definition of the Rokhlin dimension for the actions of residually compact groups, it is motivated by the definitions for actions of residually finite groups and the reals as presented in [SWZ19, Section 5] and [HSWW17, Definition 5.1], both in turn motivated by [HWZ15].
Definition 5.1.
@@
${\rm dim}_{\rm Rok}(\alpha)=\sup_{n\in\mathbb{N}}\{{\rm dim}_{\rm Rok}(\alpha,H)\colon H\leq G,H\ \text{is closed cocompact}\}.$
@@
${\rm dim}_{\rm Rok}^{\rm c}(\alpha)=\sup_{n\in\mathbb{N}}\{{\rm dim}_{\rm Rok}^{\rm c}(\alpha,H)\colon H\leq G,H\ \text{is closed cocompact}\}.$
@@ 7. $D$-absorption of crossed products
Since $A$ is $D$-absorbing, there exists a unital $\ast$-homomorphism $\tilde{k}\colon D\to F_{\infty}(A)$. As $D$ is strongly self-absorbing, $D$ is unclear. Then we can apply the Choi-Effros lifting theorem([CE76]) and find a c.p.c. lift of this map to $\ell^{\infty}(\mathbb{N},A)$, and represent it by a sequence of c.p.c. maps $k_{n}\colon D\to A$ such that $\tilde{k}(d)$ is the image of $(k_{1}(d),\ldots,k_{n}(d),\ldots)$ for all $d\in D$. Thus, we have

### v2

@@ 1. Introduction
The Rokhlin dimension was initially defined for actions by finite groups and the group $\mathbb{Z}$ of integers [HWZ15] and subsequently generalized to actions of more general groups: compact groups in [Gar17b, Gar17a, GHS21], residually finite groups in [SWZ19], and the group $\mathbb{R}$ of real numbers in [HSWW17]. All of these definitions can be used to transfer good structure properties, such as finite nuclear dimension and $\mathcal{Z}$-absorption (the latter case requires an extra assumption of commuting towers), from a C*-algebra $A$ to its crossed product C*-algebra $A\rtimes_{\alpha}G$. A natural next target would be residually compact groups, a concept that extends residual finiteness to the scope of locally compact groups. To this end, Szabó [Sza19] generalized the theory of Rokhlin dimension with commuting towers to work for second countable locally compact groups with residually compact approximations that are decreasing, discrete and normal.
In this paper, we unify all the known constructions of Rokhlin dimension and generalize the theory to work for residually compact groups, defined as follows: a second countable locally compact group is called residually compact if it admits a sequence, called a residually compact approximation, that consists of closed cocompact subgroups such that the intersection of these subgroups is trivial. Our definition of a residually compact approximation is more general than those of Szabó [Sza19] and Amini [Ami24], in that our sequence of cocompact closed subgroups need not be decreasing, nor discrete, nor normal. Unlike in the residually finite setting, residual compactness for second countable locally compact groups does not automatically admit decreasing, discrete, or normal approximating subgroups, so it is desirable to forgo these requirements in the theory. On the other hand, just like the case of residually finite groups, in order to apply the theory of Rokhlin dimension, we shall almost always assume a regularity condition on the residually compact approximations. The regular residually compact approximation in this paper includes, as a special case, the one in [Sza19] which consists of decreasing discrete normal cocompact subgroups (see Proposition 4.7). Note that the various classes of groups mentioned in the last paragraph, namely compact groups, residually finite groups and the group $\mathbb{R}$ of real numbers, are all residually compact groups. Further natural examples, for which a Rokhlin dimension theory has not been fully established, include $\mathbb{R}^{d}$ and the continuous Heisenberg group, etc. In fact, all these groups admit regular residually compact approximations.
@@
This definition of Rokhlin dimension is compatible with those given earlier in various special cases ([HWZ15, Gar17b, Gar17a, GHS21, SWZ19, HSWW17]). Imposing the extra requirement that the order zero maps $\varphi_{0},\ldots,\varphi_{d}$ have commuting ranges, we arrive at the related notion of Rokhlin dimension with commuting towers, which again generalizes the corresponding definitions in these previous papers as well as the one in [Sza19].
To see how this definition matches the description given earlier that uses Rokhlin towers, we observe that each map $\varphi_{l}$ in the above definition can be lifted to an approximately equivariant, approximately central, approximately order zero, completely positive contractive map $\psi_{l}\colon C(G/H)\to A$ (see Proposition 5.3). In particular, if $G$ is a finite group and $H$ is chosen to be trivial, then $\{\psi_{l}(\delta_{g})\colon g\in G\}$ forms a Rokhlin tower in the previous descriptions, where each $\delta_{g}$ is the characteristic function of $\{g\}$.
With the above definition, we are able to unify and generalize the previous theories on Rokhlin dimension. In particular, we show a number of important C*-algebraic properties are preserved under taking crossed products with regard to actions with finite Rokhlin dimension. Let us start with finite nuclear dimension.
@@
Another regularity property of great significance in the modern theory of C*-algebras is $\mathcal{Z}$-stability, also known as $\mathcal{Z}$-absorption, that is, the ability to tensorially absorb the Jiang-Su algebra $\mathcal{Z}$. This fits into a range of interesting properties called $D$-absorption, where $D$ is a strongly self-absorbing C*-algebra, e.g., $\mathcal{Z}$, $M_{p^{\infty}}$, $\mathcal{O}_{\infty}$ and $\mathcal{O}_{2}$ ([TW07]). For example, $\mathcal{O}_{\infty}$-absorption has played a pivotal role in the classification theory of purely infinite C*-algebras. Again, we extend the previous theories by showing that, under the assumptions of finite Rokhlin dimension with commuting towers and finite-dimensional box spaces, $D$-absorption is also preserved under taking crossed products (see [Gar17b, Theorem 4.4], [SWZ19, Theorem F], [HSWW17, Theorem C] and [Sza19, Corollary B]).
@@ 4. Box spaces and asymptotic dimension
In [Sza19], Szabó studied second countable locally compact groups equipped with a residually compact approximation, which consists of decreasing discrete normal cocompact subgroups. We point out that the residually compact approximation in [Sza19, Definition 2.1] is a regular residually compact approximation in this paper. Moreover, our regular residually compact approximation is more general than his.
@@ 5. Rokhlin dimension for actions of residually compact groups
In this section, we introduce the definition of the Rokhlin dimension for the actions of residually compact groups, it is motivated by the definitions for actions of residually finite groups and the reals as presented in [SWZ19, Definition 5.4] and [HSWW17, Definition 3.6 and Remark 3.10], both in turn motivated by [HWZ15]. Also, our definition follows in the same spirit as that of [Sza19, Definition 4.1].
Definition 5.1 ([HSWW17, Definition 5.1] and [Sza19, Definition 4.1]).
@@
${\rm dim}_{\rm Rok}(\alpha)=\sup\{{\rm dim}_{\rm Rok}(\alpha,H)\colon H\leq G,H\ \text{is closed cocompact}\}.$
@@
${\rm dim}_{\rm Rok}^{\rm c}(\alpha)=\sup\{{\rm dim}_{\rm Rok}^{\rm c}(\alpha,H)\colon H\leq G,H\ \text{is closed cocompact}\}.$
@@ 7. $D$-absorption of crossed products
Since $A$ is $D$-absorbing, there exists a unital $\ast$-homomorphism $\tilde{k}\colon D\to F_{\infty}(A)$. As $D$ is strongly self-absorbing, $D$ is nuclear. Then we can apply the Choi-Effros lifting theorem([CE76]) and find a c.p.c. lift of this map to $\ell^{\infty}(\mathbb{N},A)$, and represent it by a sequence of c.p.c. maps $k_{n}\colon D\to A$ such that $\tilde{k}(d)$ is the image of $(k_{1}(d),\ldots,k_{n}(d),\ldots)$ for all $d\in D$. Thus, we have